# Modeling zero-inflated precipitation extremes

### v1

@@ 5.1 Study Area
Khyber Pakhtunkhwa (KP) is Pakistan’s third-largest province, covering 101,741 km² area, which is $12.5\%$ of the nation’s total land. KP is situated in the northwest and shares borders with Afghanistan to the north and west, Gilgit-Baltistan to the northeast, Azad Jammu and Kashmir to the east, and the eastern regions of Punjab and Balochistan. The province is located between longitudes 69°E to 74°E and latitudes 31°N to 36°N. The temperature and precipitation patterns vary due to its diverse geography, including plains and river basins in the south and the high Hindu Kush mountains in the north. The northern region has a lot of winter precipitation, particularly in the area near the Hindu Kush. On the other hand, Peshawar and the rest of the southern region experience hot summers and comparatively dry winters due to moderate to low rainfall. Occasionally, monsoon rains affect northern areas, while summers are generally dry. This variability in precipitation significantly impacts the province’s agriculture, water resources, and climate resilience. We consider the daily average precipitation data from 2015 to 2024 for four different stations: Peshawar, Mardan, Bannu, and Hangu. For the spatial locations of considered stations, see the left panel of Figure 4. Data are downloaded from the Power NASA website https://power.larc.nasa.gov/data-access-viewer/. Every third value was retained from the data to overcome the temporal dependency. Later, we extracted the data for November, December, January, and February (the winter season), which contained more dry events. In our proposal, we are particularly interested in modeling excessive dry events (zero inflation) along with the low, moderate, and extreme wet events (see Figure 4, right panel).
@@
The ZIEGPD ($M_{1},M_{2}$ and $M_{3}$) models were separately fitted to zero-inflated rainfall intensities of each using both ML and Bayesian methods. Model $M_{1}$ provided the best fit across both approaches. Model $M_{3}$ also performed well and was comparable to model $M_{1}$, but its skewness parameter $\delta$ was sometimes poorly estimated in MLE. Model $M_{2}$ lacked flexibility in the lower tail, consistently making it the weakest performer. Model $M_{2}$ appears worst because it lacks flexibility at the bulk part. Therefore, the QQ-plot and CDF-plot of models $M_{1}$ and $M_{3}$ for stations (Peshawar, Mardan, and Bannu) are presented in Figure 5 and Figure 6, while the results of Hangu station are given in Figure S.1. As noted, Model $M_{2}$ performed the worst for all stations; the results of model $M_{2}$ are also included in SM (see e.g., Figures S.2, S.3).
@@
Figure 5: QQ-plots of fitted model to stations Peshawar, Mardan and Bannu: Model $M_{1}$ in (Top panel) and Model $M_{3}$ in (bottom panel).
@@
Table 4 presents the estimated return levels for different return periods (5, 10, 20, 50, 100, 200, and 500 years) at four rainfall stations using MLE and Bayesian estimation methods applied to two models, $M_{1}$ and $M_{3}$. The return period, also known as the average recurrence interval ($T$ years), is defined as $T=\frac{1}{P}$, where $P$ is the probability of a rainfall event occurring in a given year. It represents the expected time interval between occurrences of an extreme rainfall event of a particular magnitude or greater. The corresponding return level ($X_{T}$) represents the amount of rainfall expected to be exceeded once, on average, in that return period. For example, at the Peshawar station, using the MLE method with Model $M_{1}$, the 50-year return level is 16.2657 mm, meaning there is a 2% probability $\left(\frac{1}{50}\right)$ in any given year that rainfall will exceed this amount. The 100-year return level is 24.0664 mm, with a 1% probability (1/100) of exceedance per year. Both models show almost the same results for lower return levels, but the Bayesian method provides noticeably higher return level estimates compared to MLE for higher return periods, particularly for the 200- and 500-year return levels.
@@ 6 Conclusion
Station Method Model $\kappa$ $\delta$ $\sigma$ $\xi$ $\pi$
Peshawar MLE $M_{1}$ 0.4568 (0.3955, 0.5560) - 4.9095 (2.2781, 8.0999) 0.3281 (6.95×$\times 10^{-6}$, 0.8585) 0.5999 (0.5475, 0.6475)
$M_{3}$ 0.7645 (0.3968, 0.8288) 92.2779 (0.0005, 190.9993) 5.8624 (3.3922, 9.3927) 0.2608 (1.28×$\times 10^{-6}$, 0.5403) 0.6094 (0.5522, 0.6486)
$M_{3}$ 0.6625 (0.5079, 0.9013) 86.8341 (24.3858, 184.0118) 7.0006 (3.0222, 11.7988) 0.3165 (1.18×$\times 10^{-6}$, 0.6462) 0.6000 (0.5475, 0.6445)
Mardan MLE $M_{1}$ 0.3989 (0.3452, 0.4980) - 5.7098 (2.3001, 10.3763) 0.3703 (2.6×$\times 10^{-6}$, 0.9899) 0.6075 (0.5574, 0.6575)
$M_{3}$ 0.3981 (0.3480, 0.7329) 1.2852 (2.6×$\times 10^{-5}$, 0.9999) 6.1497 (2.9834, 11.3623) 0.35681 (1.4×$\times 10^{-6}$, 0.8340) 0.6075 (0.5550, 0.6558)
Bayesian $M_{1}$ 0.3990 (0.2987, 0.4561) - 5.7087 (4.0796, 11.7988) 0.3703 (0.21×$\times 10^{-6}$, 0.6788) 0.6075 (0.5599, 0.6540)
$M_{3}$ 0.5648 (0.4579, 0.7141) 117.3311 (43.6848, 245.4271) 9.8552 (5.2593, 14.8797) 0.3494 (1.02×$\times 10^{-6}$, 0.5547) 0.6075 (0.5621, 0.6505)
Bannu MLE $M_{1}$ 0.4772 (0.4032, 0.6304) - 3.6156 (1.3766, 5.9828) 0.2681 (1.34×$\times 10^{-6}$, 0.8212) 0.6450 (0.5964, 0.6898)
$M_{3}$ 0.4773 (0.4070, 0.6936) 0.8852 (4.12×$\times 10^{-6}$, 20.4997) 3.5023 (0.9387, 6.5650) 0.2730 (1.20×$\times 10^{-6}$, 0.7446) 0.6450 (0.5998, 0.6923)
Bayesian $M_{1}$ 0.4772 (0.3668, 0.6168) - 3.6129 (1.9140, 6.3968) 0.2686 (0.12×$\times 10^{-6}$, 0.8046) 0.6450 (0.5962, 0.6880)
$M_{3}$ 0.6487 (0.4639, 0.8189) 58.7328 (10.2062, 170.7158) 6.1288 (4.4593, 8.7845) 0.2720 (3.28×$\times 10^{-6}$, 0.8412) 0.6450 (0.5998, 0.6865)
$M_{3}$ 0.4426 (0.3899, 0.8688) 1.1696 (1.08×$\times 10^{-7}$, 98.8499) 4.1883 (0.4448, 8.0281) 0.3737 (2.92×$\times 10^{-6}$, 0.9998) 0.6100 (0.5631, 0.6551)
$M_{3}$ 0.5996 (0.4217, 0.8781) 96.1587 (26.8966, 194.4160) 7.9169 (2.8348, 9.7571) 0.3585 (1.01×$\times 10^{-6}$, 0.7981) 0.6100 (0.5643, 0.6528)
Table 4: Future return levels for the stations (Peshawar, Mardan, Bannu, and Hangu).
Station Method Model 5 (0.800) 10 (0.900) 20 (0.950) 50 (0.980) 100 (0.990) 200 (0.995) 500 (0.998)
Peshawar MLE $M_{1}$ 1.2664 4.2433 8.5123 16.2657 24.0464 33.8878 50.9241
$M_{3}$ 1.0612 4.0101 8.3478 16.0180 23.3941 32.3345 47.0180
Bayesian $M_{1}$ 1.2662 4.2423 8.5100 16.2616 24.0410 33.8816 50.9180
$M_{3}$ 1.0250 4.2514 9.3581 18.9238 28.5801 40.7636 61.7175
Mardan MLE $M_{1}$ 1.0843 4.2036 8.9872 18.0961 27.5912 39.9774 62.1878
$M_{3}$ 1.0865 4.2101 8.9839 18.0346 27.4223 39.6008 61.2640
Bayesian $M_{1}$ 1.0841 4.2028 8.9854 18.0925 27.5855 39.9689 62.1738
$M_{3}$ 0.9227 4.7319 11.3624 24.4385 38.1035 55.7957 87.0995
Bannu MLE $M_{1}$ 0.7189 2.7532 5.6242 10.6295 15.4322 21.2657 30.8880
$M_{3}$ 0.7178 2.7508 5.6232 10.6374 15.4557 21.3175 31.0104
Bayesian $M_{1}$ 0.7185 2.7518 5.6223 10.6284 15.4338 21.2724 30.9072
$M_{3}$ 0.7125 2.7298 5.6050 10.6720 15.5739 21.5617 31.4971
Hangu MLE $M_{1}$ 0.9170 3.3016 6.8749 13.6634 20.7733 30.1077 46.9903
$M_{3}$ 0.9190 3.3097 6.8833 13.6507 20.7143 29.9547 46.5841
Bayesian $M_{1}$ 0.9165 3.3005 6.8734 13.6622 20.7735 30.1110 47.0019
$M_{3}$ 0.9084 3.2791 6.8674 13.7360 20.9440 30.3887 47.3798
@@
By using CDF given in (1) and $\mathcal{W}(u;\Psi)=u^{\kappa}$, $\Psi=\kappa>0$ for (7), we found that the density function of ZIEGP distribution as
@@
By using CDF given in (1) and $\mathcal{W}(u;\Psi)=[1-B_{\delta}\{(1-u)^{\delta}\}]^{\kappa/2}$, $\Psi=(\delta,\kappa)$ with $\delta>0$ and $\kappa>0$. for (7), we found that the density function of ZIEGP distribution as
@@
Touqeer Ahmad acknowledges support from the Règion Bretagne through project SAD-2021- MaEVa.
@@
Figure S.1: QQ-plots of fitted models ($M_{1}$ and $M_{3}$) to stations Hangu (top panel), CDF plots (bottom panel).
@@
Figure S.2: QQ-plots plots of fitted model $M_{2}$ to staions Peshawar, Mardan Bannu and Hangu.

### v2

@@ 5.1 Study Area
Khyber Pakhtunkhwa (KP) is Pakistan’s third-largest province, covering 101,741 km² area, which is $12.5\%$ of the nation’s total land. KP is situated in the northwest and shares borders with Afghanistan to the north and west, Gilgit-Baltistan to the northeast, Azad Jammu and Kashmir to the east, and the eastern regions of Punjab and Balochistan. The province is located between longitudes 69°E to 74°E and latitudes 31°N to 36°N. The temperature and precipitation patterns vary due to its diverse geography, including plains and river basins in the south and the high Hindu Kush mountains in the north. The northern region has a lot of winter precipitation, particularly in the area near the Hindu Kush. On the other hand, Peshawar and the rest of the southern region experience hot summers and comparatively dry winters due to moderate to low rainfall. Occasionally, monsoon rains affect northern areas, while summers are generally dry. This variability in precipitation has a significant impact on the province’s agriculture, water resources, and climate resilience. We consider the daily average precipitation data from 2015 to 2024 for four different stations: Peshawar, Mardan, Bannu, and Hangu. For the spatial locations of the considered stations, see the left panel of Figure 4. Data are downloaded from the Power NASA website https://power.larc.nasa.gov/data-access-viewer/. Every third value was retained from the data to overcome the temporal dependency. Later, we extracted the data for November, December, January, and February (the winter season), which contained more dry events. We additionally replaced the values below $0.1mm$ by zero, considering it a dry day event. In our proposal, we are particularly interested in modeling excessive dry events (zero inflation) along with the low, moderate, and extreme wet events (see Figure 4, right panel).
@@
The ZIEGPD ($M_{1},M_{2}$ and $M_{3}$) models were separately fitted to zero-inflated rainfall intensities of each using both ML and Bayesian methods. Model $M_{1}$ provided the best fit across both approaches. Model $M_{3}$ also performed well and was comparable to model $M_{1}$, but its skewness parameter $\delta$ was sometimes poorly estimated in MLE. Model $M_{2}$ lacked flexibility in the lower tail, consistently making it the weakest performer. Therefore, the QQ-plot and CDF-plot of models $M_{1}$ and $M_{3}$ for stations (Peshawar, Mardan, and Bannu) are presented in Figure 5 and Figure 6, while the results of Hangu station are given in Figure S.2. As noted, Model $M_{2}$ performed the worst for all stations; the results of model $M_{2}$ are also included in SM (see e.g., Figures S.3, S.4). As observed in both the simulation and real-data analyses, the model $M_{2}$ exhibits poor performance, primarily due to its limited flexibility in capturing the lower tail, even in the presence of a substantial proportion of zeros. We nevertheless included this model in our study to serve as a benchmark: despite incorporating a zero-inflated component, its shortcomings remain evident, similar to Naveau et al., (2016), thereby highlighting the benefits of more flexible alternatives, for instance, models $M_{2}$ and $M_{3}$.
@@
Figure 5: $Q$-plots of fitted model to stations Peshawar, Mardan and Bannu: Model $M_{1}$ in (Top panel) and Model $M_{3}$ in (bottom panel).
@@
Station Method Model $\kappa$ $\delta$ $\sigma$ $\xi$ $\pi$
[…]
Table 4 presents the estimated return levels for different return periods (5, 10, 20 years) at four rainfall stations using MLE and Bayesian estimation methods applied to two models, $M_{1}$ and $M_{3}$. The return period, also known as the average recurrence interval ($T$ years), is defined as $T=\frac{1}{P}$, where $P$ is the probability of a rainfall event occurring in any given year. It represents the expected time interval between occurrences of an extreme rainfall event of a particular magnitude or greater. The corresponding return level ($X_{T}$) represents the amount of rainfall expected to be exceeded once, on average, in that return period.
For example, at the Peshawar station, the MLE method with models $M_{1}$ (resp. $M_{3}$) estimates a 10-year return level of 4.2433 mm (resp. 4.0101 mm), implying that in any given year there is a 10% probability (1/10) that rainfall will exceed this amount. Similarly, the 20-year return level is 8.5123 mm for $M_{1}$ and 8.3478 mm for $M_{3}$, which correspond to a 5% annual exceedance probability (1/20). At shorter return periods (e.g., 5 years), both $M_{1}$ and $M_{3}$ provide very similar estimates, indicating that the choice of model has a limited influence on lower return levels. However, as the return period increases, differences between the models become more evident, reflecting greater sensitivity of extreme rainfall estimates to model specification at higher quantiles.
Table 4: Return levels for the stations (Peshawar, Mardan, Bannu, and Hangu).
Station Method Model 5 (0.800) 10 (0.900) 20 (0.950)
Peshawar MLE $M_{1}$ 1.2664 4.2433 8.5123
$M_{3}$ 1.0612 4.0101 8.3478
Bayesian $M_{1}$ 1.2662 4.2423 8.5100
$M_{3}$ 1.0250 4.2514 9.3581
Mardan MLE $M_{1}$ 1.0843 4.2036 8.9872
$M_{3}$ 1.0865 4.2101 8.9839
Bayesian $M_{1}$ 1.0841 4.2028 8.9854
$M_{3}$ 0.9227 4.7319 11.3624
Bannu MLE $M_{1}$ 0.7189 2.7532 5.6242
$M_{3}$ 0.7178 2.7508 5.6232
Bayesian $M_{1}$ 0.7185 2.7518 5.6223
$M_{3}$ 0.7125 2.7298 5.6050
Hangu MLE $M_{1}$ 0.9170 3.3016 6.8749
$M_{3}$ 0.9190 3.3097 6.8833
Bayesian $M_{1}$ 0.9165 3.3005 6.8734
$M_{3}$ 0.9084 3.2791 6.8674
@@ 6 Conclusion
Appendix A. By using CDF given in (1) and $\mathcal{W}(u;\Psi)=u^{\kappa}$, $\Psi=\kappa>0$ for (7), we found that the density function of ZIEGP distribution as
@@
By using CDF given in (1) and $\mathcal{W}(u;\Psi)=[1-B_{\delta}\{(1-u)^{\delta}\}]^{\kappa/2}$, $\Psi=(\delta,\kappa)$ with $\delta>0$ and $\kappa>0$. For (7), we found that the density function of the ZIEGP distribution as
@@
Appendix B. In case of non-stationary data, we also developed the non-stationary version of the proposed models. It is a semi-parametric regression model that assumes a parametric distribution for the response variable, with the option for the distribution parameters to change according to explanatory variables or random effects. In non-stationary framework, the observation $z_{t},t=1,\cdots,n$ and fitted to distribution $h(z_{t}|\theta^{t})$ with $\theta^{t}=(\theta_{1t},\cdots,\theta_{pt})$ where $p$ is the number of distribution parameters at time $t$. To relate the parameters of the distribution to time $t$ or vector of covariates $(\boldsymbol{x_{i}})$ of dimension $p$, we consider a generic structure of additive predictors of the form
$\eta_{j}^{\theta^{t}}=\beta_{j0}^{{\theta^{t}}}+f_{j1}^{{\theta^{t}}}(\boldsymbol{x}_{ij1})+\dots+f_{jk}^{{\theta^{t}}}(\boldsymbol{x}_{ijk})+\dots+f_{jK_{j}}^{{\theta^{t}}}(\boldsymbol{x}_{ijK_{j}}),$ (A.4)
@@
Touqeer Ahmad acknowledges support from the Project JUNON.
@@
S.1 Simulation Study
((a))
((b))
((c))
Figure S.1: Box plots showing the maximum likelihood estimates and Bayesian estimates of model parameters for each model type, based on sample sizes of $n=500$ with $10^{4}$ replications under varying parameter settings.
@@
Figure S.2: $Q$-plots of fitted models ($M_{1}$ and $M_{3}$) to stations Hangu (top panel), CDF plots (bottom panel).
@@
Figure S.3: $Q$-plots plots of fitted model $M_{2}$ to staions Peshawar, Mardan Bannu and Hangu.
@@
Algorithm 1 Nonparametric Bootstrap for ZIEGPD Confidence Intervals
1:Input: Observed data $\{z_{1},\ldots,z_{n}\}$, number of bootstrap replicates $B$, confidence level $(1-\alpha)$.
2:Fit the ZIEGPD model to $\{z_{1},\ldots,z_{n}\}$ and obtain parameter estimates
[…]
for example, under Model $M_{1}$.
3:Initialize bootstrap storage
$\{\pi^{*(b)},\kappa^{*(b)},\sigma^{*(b)},\xi^{*(b)}\}_{b=1}^{B}.$
4:for $b=1,\ldots,B$ do
5:  Draw a bootstrap resample $\{z_{1}^{*},\ldots,z_{n}^{*}\}$ from $\{z_{1},\ldots,z_{n}\}$ with replacement.
6:  Fit the ZIEGPD model to $\{z_{1}^{*},\ldots,z_{n}^{*}\}$ and obtain
$\hat{\theta}^{*(b)}=\big(\pi^{*(b)},\kappa^{*(b)},\sigma^{*(b)},\xi^{*(b)}\big).$
7:  Store $\hat{\theta}^{*(b)}$.
8:end for
9:Construct percentile confidence intervals for each parameter $\theta_{j}\in\{\pi,\kappa,\sigma,\xi\}$:
$CI_{\theta_{j}}=\Big[Q_{\alpha/2}\big(\{\theta_{j}^{*(b)}\}_{b=1}^{B}\big),\;Q_{1-\alpha/2}\big(\{\theta_{j}^{*(b)}\}_{b=1}^{B}\big)\Big],$
where $Q_{q}(\cdot)$ denotes the empirical $q$-th quantile.
10:return $\hat{\theta}$ and the confidence intervals $\{CI_{\pi},CI_{\kappa},CI_{\sigma},CI_{\xi}\}$.